OpenStudy (jpharper12345678cars):

What is the value of the function when x = –1? A. –8 B. –4 C. 2 D. 6

OpenStudy (jpharper12345678cars):

i only have five qestions

OpenStudy (jpharper12345678cars):

and i will fan and medal

OpenStudy (john_es):

But you should write the formula of the function.

OpenStudy (jpharper12345678cars):

i suck at math

OpenStudy (john_es):

Then, you only need to improve. But first, you should give us the formula of the function.

OpenStudy (jpharper12345678cars):

oh sorry here y=3x-1 x= -1

OpenStudy (john_es):

Well, the problem says, x=-1, so you should find the value of y when x=-1.

OpenStudy (jpharper12345678cars):

-4

OpenStudy (john_es):

Yes

OpenStudy (jpharper12345678cars):

can you help me with four more please

OpenStudy (john_es):

Ok, tell me

OpenStudy (jpharper12345678cars):

What is the value of the function v =3/2t - 5/2 when t = 3?

OpenStudy (john_es):

Same as before, put a 3 where the t is, and calculate the result.

OpenStudy (john_es):

Do you know how to sum these numbers?

OpenStudy (jpharper12345678cars):

is it 2?

OpenStudy (john_es):

Perfect!

OpenStudy (jpharper12345678cars):

Evaluate the function y = 20 – x when x = –3 A. 17 B. 60 C. 23 D. –60

OpenStudy (jpharper12345678cars):

20- (-3) = 23?

OpenStudy (john_es):

Perfect!

OpenStudy (jpharper12345678cars):

Find the value of the function when n = 6. p = 0.008n A. 0.048 B. 0.0042 C. 0.48 D. 6.008

OpenStudy (john_es):

So, what do you think this time?

OpenStudy (jpharper12345678cars):

p=0.008 * 6= 0.048 so p=0.048

OpenStudy (john_es):

Yeah, perfect!

OpenStudy (jpharper12345678cars):

What is the value of the function when x = 6? y = 4x + 5

OpenStudy (jpharper12345678cars):

4*6+5=29?

OpenStudy (john_es):

Perfect!

OpenStudy (jpharper12345678cars):

can you please oh please help me with more you got me to pass thank you thank you thank you

OpenStudy (john_es):

Ok, tell me

OpenStudy (jpharper12345678cars):

you sure im not being a bother

OpenStudy (john_es):

No problem, tell me

OpenStudy (jpharper12345678cars):

Evaluate the function at x = –2. A. y = –4 B. y = 0 C. y = 2 D. y = 3

OpenStudy (jpharper12345678cars):

c?

OpenStudy (john_es):

I suppose these are two questions. The first one, the value of the function at x=-2, you can tell me the answer, but you should give me again the function if you want me to check it

OpenStudy (jpharper12345678cars):

no d?

OpenStudy (jpharper12345678cars):

no the link is the gragh

OpenStudy (john_es):

Ok, now I understand

OpenStudy (jpharper12345678cars):

i think its d?

OpenStudy (john_es):

You are right

OpenStudy (john_es):

It is D

OpenStudy (jpharper12345678cars):

What is the value of the function when x = 2? x −4 −2 0 2 4 y 3 4 2 6 7

OpenStudy (jpharper12345678cars):

im lost on this one

OpenStudy (jpharper12345678cars):

+2

OpenStudy (john_es):

You should find the expression of the function. You know that, in general, a linear function is of the form\[y=mx+n\] where you can find the slope, and the intersection n. The intersection n can be find because when x=0, then\[2=n\]. You can find the slope, using,\[m=\frac{y_2-y_1}{x_2-x_1}=\frac{6-2}{2-0}=2\]. So the function is,\[y=2x+2\] And when x=2, then y is\[y=2\cdot2+2=6\]

OpenStudy (jpharper12345678cars):

omg thank you so very very very much

OpenStudy (john_es):

But don't worry, the problem is easier. You don't need to find all of that, simply find the x in the table and you will see that when x=2, y=6. See you.